Thank you for helping me> I really appreciate you taking time out of your
schedule. Thanks again!
The revenue for a sandwich shop is directly proportional to its advertising budget. The owner spent $2,000 a month for advertising, the revenue is $120,000.If the revenue now is $180,000, how much is the owner spending on advertising?

Answer by stanbon(75887)

The revenue for a sandwich shop is directly proportional to its advertising budget. The owner spent $2,000 a month for advertising, the revenue is $120,000.If the revenue now is $180,000, how much is the owner spending on advertising?
-----------------
x/180000 = 2000/120000
x/180000 = 1/60
x = 180000/60
x = $3000 for advertising
==================
Cheers,
Stan H.